What is the answers to the folloeing promblem
y=2x
y=x-3
y=-1+2

Answer by Alan3354(69443)

What is the answers to the folloeing promblem
y=2x
y=x-3
y=-1+2
--------------
Is this one problem or 3?
What do you want to solve for?
For this one:
y=-1+2
y = 1